OpenStudy (anonymous):

verify the equation is an identity: sin (x)/1-sin ^2 (x)

OpenStudy (anonymous):

is it \[\frac{ \sin x }{ 1-\sin^2x }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Is tan x, considered an identity?

OpenStudy (anonymous):

I think so

OpenStudy (raden):

we have to know, it would be equals what ?

OpenStudy (anonymous):

sec x tan x

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

okay, that makes more sense.

OpenStudy (raden):

use these identities : 1 - sin^2 x = cos^2 x sin/cos = tan 1/cos = sec

OpenStudy (anonymous):

pretty much, should not take more than 4 steps.

OpenStudy (raden):

sin (x)/1-sin ^2 (x) = sin (x)/cos^2 x = sin(x)/cos(x) * 1/cos(x) = tan(x) sec(x)

OpenStudy (anonymous):

or sec x tan x 1/cos x sin x /cos x sinx / cos 2x sinx / 1-sin2x

OpenStudy (anonymous):

what we have here is simply a simplification trig problem... wherein we simplify... \[\frac{ \sin x }{ 1-\sin^2x }=\sec x \tan x\]

OpenStudy (anonymous):

I mean cosx squared and sin x squared not double angle.

OpenStudy (anonymous):

using trig identity of course... :)